OpenStudy (luigi0210):

Integrate using a trig sub:

OpenStudy (luigi0210):

\[\LARGE \frac{x^3}{\sqrt{1-x^2}}~dx\]

OpenStudy (abb0t):

use trig sub :)

OpenStudy (abb0t):

oh, and parts.

OpenStudy (luigi0210):

I just noticed I forgot my integral.. I must be tired \[\LARGE \int \frac{x^3}{\sqrt{1-x^2}}~dx\] The \(\Large \int \frac{1}{\sqrt{1-x^2}}dx=sin^{-1}x\)?

OpenStudy (abb0t):

I forgot all my math, but I think that is correct.

OpenStudy (anonymous):

sub \(x=\sin \theta\) \(dx= \cos \theta d\theta\) \(\large \int \frac {(\sin \theta)^3 \cos \theta d\theta }{ \sqrt {1-\sin^2 \theta }}\) \(\large \int \frac {(\sin \theta)^3 \cos \theta d\theta }{ \cos \theta }\) \(\large \int (\sin \theta)^3 d \theta \)